# Advancing Pose-Guided Image Synthesis with Progressive Conditional Diffusion Models

### v1

@@ Abstract
Recent work has showcased the significant potential of diffusion models in pose-guided person image synthesis. However, owing to the inconsistency in pose between the source and target images, synthesizing an image with a distinct pose, relying exclusively on the source image and target pose information, remains a formidable challenge. This paper presents Progressive Conditional Diffusion Models (PCDMs) that incrementally bridge the gap between person images under the target and source poses through three stages. Specifically, in the first stage, we design a simple prior conditional diffusion model that predicts the global features of the target image by mining the global alignment relationship between pose coordinates and image appearance. Then, the second stage establishes a dense correspondence between the source and target images using the global features from the previous stage, and an inpainting conditional diffusion model is proposed to further align and enhance the contextual features, generating a coarse-grained person image. In the third stage, we propose a refining conditional diffusion model to utilize the coarsely generated image from the previous stage as a condition, achieving texture restoration and enhancing fine-detail consistency. The three-stage PCDMs work progressively to generate the final high-quality and high-fidelity synthesized image. Both qualitative and quantitative results demonstrate the consistency and photorealism of our proposed PCDMs under challenging scenarios.The code and model will be available at https://github.com/muzishen/PCDMs.
@@ 1 Introduction
Given an image of a specific person under a particular pose, pose-guided image synthesis (Zhang et al., 2022; Ren et al., 2022; Bhunia et al., 2023) aims to generate images of the person with the same appearance and meanwhile under the given target pose, which more importantly, are expected to be as photorealistic as possible. It holds broad and robust application potential in e-commerce and content generation. Meanwhile, the generated images can be used to improve the performance of downstream tasks, such as person re-identification (Ye et al., 2021). However, since pose disparities between the source and target images, generating an image with a different pose solely based on the source image and target pose information remains a significant challenge.
@@
Figure 1: (a) Existing methods typically utilize unaligned image-to-image generation at the conditional level. (b) Our approach progressively predicts the global features, dense correspondences, and texture restoration of target image, enabling image synthesis.
@@ 2 Related Work
Person Image Synthesis. The task of person image synthesis has achieved great development during these years, especially with the unprecedented success of deep learning. The earlier methods (Ma et al., 2017; Men et al., 2020) treat the synthesis task as conditional image generation, using conditional generative adversarial networks (CGANs) (Mirza & Osindero, 2014) to generate the target image with the source appearance image and target pose as conditions. However, due to the inconsistency between the source and target poses, the effectiveness of directly connecting the source image with the target pose is limited. To overcome this challenge, VUnet (Esser et al., 2018) adopts a joint application of VAE and U-Net to decouple the appearance and pose of the character image. Furthermore, Def-GAN (Siarohin et al., 2018) proposes a deformable GAN that decomposes the overall deformation through a set of local affine transformations to address the misalignment issues caused by different poses. On the other hand, some works (Liu et al., 2019; Ren et al., 2020) utilize flow-based deformation to transform source information, improving pose alignment. For example, GFLA (Ren et al., 2020) obtains global flow fields and occlusion masks to warp local patches of the source image to match the desired pose. Similarly, ClothFlow (Han et al., 2019) is a model designed based on this concept for segmentation-guided human image generation. Subsequently, methods such as PISE (Zhang et al., 2021), SPGnet (Lv et al., 2021), and CASD (Zhou et al., 2022) leverage parsing maps to generate the final image. CoCosNet (Zhou et al., 2021) extracts dense correspondence between cross-domain images through attention-based operations. However, person image synthesis is essentially a transformation from non-aligned images to images, and the absence of global appearance features of the target image can lead to less realistic results.
Figure 2: The three-stage pipeline of Progressive Conditional Diffusion Models (PCDMs) progressively operates to generate the final high-quality and high-fidelity synthesized image.
@@ 3.2 Prior conditional diffusion model
As depicted in Figure 3, the prior conditional diffusion model is a transformer network, conditioned on the pose of the source image, the pose of the target image, and the source image. We first adopt OpenPose (Cao et al., 2017) to acquire the pose coordinates for the pose of source and target images. A compact trainable pose network, composed of 3 linear layers, is used to project the pose coordinates into the pose embedding. For the source image, we also use a CLIP image encoder to extract the image embedding and add a linear layer to project the image embedding. Additionally, we add an extra embedding to predict the target global embedding. The above embeddings plus timestep embedding and noisy image embedding of the target image are concatenated into a sequence of embeddings as the input of the transformer network.
@@
$\vspace{-5pt}L^{\mathrm{prior}}=\mathbb{E}_{\bm{x}_{0},\bm{x}_{s},\bm{p}_{s},\bm{p}_{t},t}\|\bm{x}_{0}-\bm{x}_{\theta}\big{(}\bm{x}_{t},\bm{x}_{s},\bm{p}_{s},\bm{p}_{t},t\big{)}\|^{2}.\vspace{-5pt}$ (3)
@@ 3.4 Refining conditional diffusion model
$\vspace{-5pt}L^{\mathrm{prior}}=\mathbb{E}_{\bm{x}_{0},\bm{\epsilon},\bm{i}_{ct},\bm{x}_{s},t}\|\bm{\epsilon}-\bm{\epsilon}_{\theta}\big{(}\bm{x}_{t},\bm{i}_{ct},\bm{x}_{s},t\big{)}\|^{2}.\vspace{-5pt}$ (7)
@@ 4 Experiments
Metrics. We conduct a comprehensive evaluation of the model, considering both objective and subjective metrics. Objective indicators include structural similarity index measure (SSIM) (Wang et al., 2004), learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018), and fréchet inception distance (FID) (Heusel et al., 2017). In contrast, subjective assessments prioritize user-oriented metrics (Zhou et al., 2022), including the percentage of real images misclassified as generated images (R2G), the percentage of generated images misclassified as real images (G2R), and the percentage of images deemed superior among all models (Jab).
@@ 4.1 Quantitative and Qualitative Results
Quantitative Results. As shown in Table 1, our approach excels in two out of three metrics on the DeepFashion compared to other models, regardless of whether it is based on GAN, VAE, flow-based model, or diffusion model. For example, when using the flow framework (i.e., CASD), PCDMs outperforms the SOTA flow-based method, CASD, even without explicitly decoupling pose and style. Although PCDMs have a lower FID score than PIDM (which also employs the diffusion model), we surpass it on the other two metrics, and subsequent experiments further demonstrate our method’s superiority.
@@
Qualitative Results. As shown in Figure 6, we comprehensively compare our PCDMs and other state-of-the-art methods on the DeepFashion dataset. Several observations can be drawn from the results: (1) despite the minuscule size of the necklace in the source image (as seen in the first and second rows), only our proposed PCDMs and the PIDM, which also utilizes a diffusion model, can focus on it. But PCDMs can generate higher-quality images compared to the PIDM. (2) In scenarios involving extreme poses and large area occlusions (as seen in the third and fourth rows), only our method can generate images that align reasonably with the target. This can be attributed to our method’s ability to capture and enhance the global features of the target image. (3) In situations with complex textures and numbers (as seen in the fifth to seventh rows), our method significantly surpasses others in preserving appearance textures, primarily due to our method’s capability to refine textures and maintain consistency. To sum up, our method consistently produces more realistic and lifelike person images, demonstrating the advantage of our PCDMs’ multi-stage progressive generation approach. See C.1 for more examples.
@@
This study included comparisons with fundamental facts (namely, R2G and G2R) and comparisons with other methods (i.e., J2b). A higher score on these three metrics indicates superior performance. As shown in Figure 7, PCDMs exhibit commendable performance across all three baseline metrics on DeepFashion. For instance, the proportion of PCDMs images perceived as real in all instances is 56.2% (G2R), nearly 20.4% higher than the next best model. Our Jab score stands at 44.1%, suggesting a preference for our method among the participants. Please refer to C.4 of the Appendix for more detail.
Figure 8: Qualitative ablation results on the DeepFashion. See C.2 for more examples.
@@ 4.2 Ablation Study
To demonstrate the efficacy of each stage introduced within this paper, we have devised the following variations, all of which fall under the PCDMs framework but incorporate distinct configurations. B1 represents using the inpainting conditional diffusion model only. B2 denotes the simultaneous use of the prior conditional diffusion model and the inpainting conditional diffusion model without using the refining conditional diffusion model. B3 stands for the simultaneous use of the redrawn conditional diffusion model and the refining conditional diffusion model, excluding the prior conditional diffusion model.
Figure 8 shows the impact of each stage on the DeepFashion dataset. B1 can generate person images that roughly conform to the target pose. However, there are severe distortions in the generated image, such as limb loss and confusion between appearance and limbs. This indicates that generating images from unaligned inputs is a highly challenging task without global features. In contrast, in B2, once the global features of the target image obtained from the prior conditional diffusion model are added, the generated person images basically match the target pose in structure. Although the personal images generated by B2 can retain the appearance of the source image is limited, it is already very close to the real image. This shows that the refining conditional diffusion model can establish a dense correspondence between the source and target images, enhancing the contextual features. In addition, from a visual perception perspective, B3 is superior to B1 and B2 regarding detail texture, although it is slightly inferior to B2 regarding pose coordination. Finally, when we use the PCDMs of the three stages in the last column, it is visually superior to B1, B2, and B3. This indicates that when dealing with complex poses, the PCDMs of the three stages can gradually produce visually more satisfying results.
To more comprehensively validate the effectiveness of the proposed refining conditional diffusion model, we apply it to other state-of-the-art (SOTA) methods. As shown in Figure 9, the first and fourth rows denote the source image and ground truth (target image). The second and third rows present the results before and after refinement via refining the conditional diffusion model.
@@
We can observe that the refining conditional diffusion model significantly improves the results of all state-of-the-art methods. For instance, when dealing with ADGAN and PIDM, our method helps to fill in minor details, such as missing belts and hats, thereby enhancing the completeness of the generated images. For methods like GFLA, DPTN, CASD, and NTED, our model can finely process textures, maintain shape and texture consistency, and generate more explicit and realistic images. These results indicate that the refining conditional diffusion model has universality across different state-of-the-art methods, offering potential improvements for various person image generation tasks.
@@ 4.3 Application
In Table 2, we further evaluate the applicability of images generated by PCDMs in a downstream task, i.e., person re-identification. We conduct these re-identification experiments on the Market-1501 dataset, adhering to the PIDM protocol (Bhunia et al., 2023). Initially, we randomly select subsets of 20%, 40%, 60%, and 80% from the total real training set of the Market-1501 dataset, ensuring each identity is represented by at least one image. This selection process yields a new dataset. Following this, we employ BoT (Luo et al., 2019) as the base network and conduct baseline training with each subset of the data. We then generate synthetic images from the new dataset, randomly selecting those that share the same identity and pose. This synthesized data is then merged with the original dataset to train BoT. The Rank1 results are presented in Table 2. The results indicate that PCDMs significantly boost the re-identification performance compared to the baseline. Furthermore, when compared with state-of-the-art methods such as PTN (Zhu et al., 2019), GFLA (Ren et al., 2020), DPTN (Zhang et al., 2022), and PIDM (Bhunia et al., 2023), PCDMs consistently demonstrate superior performance in re-identification tasks.
@@ 5 Conclusion
This paper has demonstrated the significant potential of Progressive Conditional Diffusion Models (PCDMs) in addressing the challenges of pose-guided person image synthesis. The inherent inconsistency in pose between source and target images has been a formidable obstacle in synthesizing an image with a distinct pose, relying solely on the source image and target pose information. Our proposed PCDMs incrementally bridge this gap through a three-stage process. In the first stage, a simple prior conditional diffusion model is designed to predict the global features of the target image by mining the global alignment relationship between pose coordinates and image appearance. The second stage establishes a dense correspondence between the source and target images using the global features from the previous stage, and an inpainting conditional diffusion model is proposed to further align and enhance the contextual features, generating a coarse-grained person image. In the final stage, a refining conditional diffusion model is proposed to utilize the coarsely generated image from the previous stage as a condition, achieving texture restoration and enhancing fine-detail consistency. The three stages of the PCDMs work progressively to generate the final high-quality and high-fidelity synthesized image. Both qualitative and quantitative results have demonstrated the consistency and photorealism of our proposed PCDMs under challenging scenarios.
Futher Work. Our method significantly improves person image synthesis quality, while the use of two additional prior and refining models leads to increased computational resource consumption and longer inference time. Future work should explore efficient methods that provide equivalent or superior quality while reducing computational overhead and inference time.
@@ 6 Ethics Statement
7 Reproducibility
The data, code, and models will be publicly available.

### v3

@@ Abstract
Recent work has showcased the significant potential of diffusion models in pose-guided person image synthesis. However, owing to the inconsistency in pose between the source and target images, synthesizing an image with a distinct pose, relying exclusively on the source image and target pose information, remains a formidable challenge. This paper presents Progressive Conditional Diffusion Models (PCDMs) that incrementally bridge the gap between person images under the target and source poses through three stages. Specifically, in the first stage, we design a simple prior conditional diffusion model that predicts the global features of the target image by mining the global alignment relationship between pose coordinates and image appearance. Then, the second stage establishes a dense correspondence between the source and target images using the global features from the previous stage, and an inpainting conditional diffusion model is proposed to further align and enhance the contextual features, generating a coarse-grained person image. In the third stage, we propose a refining conditional diffusion model to utilize the coarsely generated image from the previous stage as a condition, achieving texture restoration and enhancing fine-detail consistency. The three-stage PCDMs work progressively to generate the final high-quality and high-fidelity synthesized image. Both qualitative and quantitative results demonstrate the consistency and photorealism of our proposed PCDMs under challenging scenarios. The code and model will be available at https://github.com/tencent-ailab/PCDMs.
@@ 1 Introduction
Given an image of a specific person under a particular pose, pose-guided image synthesis (Zhang et al., 2022; Ren et al., 2022; Bhunia et al., 2023) aims to generate images of the person with the same appearance and meanwhile under the given target pose, which more importantly, are expected to be as photorealistic as possible. It holds broad and robust application potential in e-commerce and content generation. Meanwhile, the generated images can be used to improve the performance of downstream tasks, such as person re-identification (Ye et al., 2021; Shen et al., 2023b). However, since pose disparities between the source and target images, generating an image with a different pose solely based on the source image and target pose information remains a significant challenge.
@@
Figure 1: Existing methods typically utilize unaligned image-to-image generation at the conditional level.
@@ 2 Related Work
Person Image Synthesis. The task of person image synthesis has achieved great development during these years, especially with the unprecedented success of deep learning. The earlier methods (Ma et al., 2017; Men et al., 2020) treat the synthesis task as conditional image generation, using conditional generative adversarial networks (CGANs) (Mirza & Osindero, 2014) to generate the target image with the source appearance image and target pose as conditions. However, due to the inconsistency between the source and target poses, the effectiveness of directly connecting the source image with the target pose is limited. To overcome this challenge, VUnet (Esser et al., 2018) adopts a joint application of VAE and U-Net to decouple the appearance and pose of the character image. Furthermore, Def-GAN (Siarohin et al., 2018) proposes a deformable GAN that decomposes the overall deformation through a set of local affine transformations to address the misalignment issues caused by different poses. On the other hand, some works (Liu et al., 2019; Ren et al., 2020) utilize flow-based deformation to transform source information, improving pose alignment. For example, GFLA (Ren et al., 2020) obtains global flow fields and occlusion masks to warp local patches of the source image to match the desired pose. Similarly, ClothFlow (Han et al., 2019) is a model designed based on this concept for segmentation-guided human image generation. Subsequently, methods such as PISE (Zhang et al., 2021), SPGnet (Lv et al., 2021), and CASD (Zhou et al., 2022) leverage parsing maps to generate the final image. CoCosNet (Zhou et al., 2021) extracts dense correspondence between cross-domain images through attention-based operations. However, person image synthesis is essentially a transformation from non-aligned images to images, and the absence of global appearance features of the target image can lead to less realistic results. Besides, there are some multi-stage methods in other generative domains. For example, Grigorev et al. (Grigorev et al., 2019) proposed a framework based on CNNs that first performs pose warping, followed by texture repair. Unselfie (LastName, 2014) introduces a pipeline that first identifies the target’s neutral pose, repairs body texture, and then perfects and synthesizes the character in the background. While these methods can fit the target pose well, they lose a sense of realism when combined with the background or the human body.
Figure 2: The three-stage pipeline of Progressive Conditional Diffusion Models (PCDMs) progressively operates to generate the final high-quality and high-fidelity synthesized image. Our approach progressively predicts the global features, dense correspondences, and texture restoration of target image, enabling image synthesis.
@@ 3.2 Prior conditional diffusion model
As depicted in Figure 3, the prior conditional diffusion model is a transformer network, conditioned on the pose of the source image, the pose of the target image, and the source image. We first adopt OpenPose (Cao et al., 2017) to acquire the pose coordinates for the pose of source and target images. A compact trainable pose network, composed of 3 linear layers, is used to project the pose coordinates into the pose embedding. For the source image, we also use a CLIP image encoder to extract the image embedding and add a linear layer to project the image embedding. Additionally, we add an extra embedding to predict the unnoised global embedding of target image. The above embeddings plus timestep embedding and noisy image embedding of the target image are concatenated into a sequence of embeddings as the input of the transformer network.
@@
$\vspace{-5pt}L^{\mathrm{prior}}=\mathbb{E}_{\bm{x}_{0},{\color[rgb]{0,0,0}{\bm{\epsilon}}},\bm{x}_{s},\bm{p}_{s},\bm{p}_{t},t}\|\bm{x}_{0}-\bm{x}_{\theta}\big{(}\bm{x}_{t},\bm{x}_{s},\bm{p}_{s},\bm{p}_{t},t\big{)}\|^{2}.\vspace{-5pt}$ (3)
@@ 3.4 Refining conditional diffusion model
$\vspace{-5pt}L^{\mathrm{refining}}=\mathbb{E}_{\bm{x}_{0},\bm{\epsilon},\bm{i}_{ct},\bm{x}_{s},t}\|\bm{\epsilon}-\bm{\epsilon}_{\theta}\big{(}\bm{x}_{t},\bm{i}_{ct},\bm{x}_{s},t\big{)}\|^{2}.\vspace{-5pt}$ (7)
@@ 4 Experiments
Metrics. We conduct a comprehensive evaluation of the model, considering both objective and subjective metrics. Objective indicators include structural similarity index measure (SSIM) (Wang et al., 2004), learned perceptual image patch similarity (LPIPS) (Zhang et al., 2018), and fréchet inception distance (FID) (Heusel et al., 2017). In contrast, subjective assessments prioritize user-oriented metrics, including the percentage of real images misclassified as generated images (R2G) (Ma et al., 2017), the percentage of generated images misclassified as real images (G2R) (Ma et al., 2017), and the percentage of images deemed superior among all models (Jab) (Siarohin et al., 2018).
@@ 4.1 Quantitative and Qualitative Results
Quantitative Results. From Table 1, PCDMs excels in two out of three metrics on the DeepFashion compared to other models, regardless of whether it is based on GAN, VAE, flow-based model, or diffusion model. For example, when using the flow framework (i.e., CASD), PCDMs outperforms the SOTA flow-based method, CASD, even without explicitly decoupling pose and style. Although PCDMs worse FID score than PIDM (which also employs the diffusion model), we surpass it on the other two metrics, and subsequent experiments further demonstrate our method’s superiority.
@@
Qualitative Results. As shown in Figure 6, we comprehensively compare our PCDMs and other state-of-the-art methods on the DeepFashion dataset. Several observations can be drawn from the results: (1) despite the minuscule size of the necklace in the source image (as seen in the first and second rows), only our proposed PCDMs and the PIDM, which also utilizes a diffusion model, can focus on it. But PCDMs can generate higher-quality images compared to the PIDM. (2) In scenarios involving extreme poses and large area occlusions (as seen in the third and fourth rows), only our method can generate images that align reasonably with the target. This can be attributed to our method’s ability to capture and enhance the global features of the target image. (3) In situations with complex textures and numbers (as seen in the fifth to seventh rows), our method significantly surpasses others in preserving appearance textures, primarily due to our method’s capability to refine textures and maintain consistency. (4) The last two rows present source images with invisible logo and target images with visible logo. The results indicate that PCDMs do not overfit, and our results demonstrate better visual consistency than other SOTA methods. To sum up, our method consistently produces more realistic and lifelike person images, demonstrating the advantage of our PCDMs’ multi-stage progressive generation approach. See C.2 for more examples.
@@
This study included comparisons with fundamental facts (namely, R2G and G2R) and comparisons with other methods (i.e., J2b). A higher score on these three metrics indicates superior performance. As shown in Figure 7, PCDMs exhibit commendable performance across all three baseline metrics on DeepFashion. For instance, the proportion of PCDMs images perceived as real in all instances is 56.2% (G2R), nearly 20.4% higher than the next best model. Our Jab score stands at 44.1%, suggesting a preference for our method among the participants. Please refer to C.6 of the Appendix for more detail.
Figure 8: Qualitative ablation results. See C.3 for more examples.
@@ 4.2 Ablation Study
To demonstrate the efficacy of each stage introduced within this paper, we have devised the following variations, all of which fall under the PCDMs framework but incorporate distinct configurations. B1 represents using the inpainting conditional diffusion model only. B2 denotes the simultaneous use of the prior conditional diffusion model and the inpainting conditional diffusion model without using the refining conditional diffusion model. B3 stands for the simultaneous use of the inpainting conditional diffusion model and the refining conditional diffusion model, excluding the prior conditional diffusion model.
Figure 8 shows the impact of each stage on the DeepFashion. B1 can generate person images that roughly conform to the target pose. However, there are severe distortions in the generated image, such as limb loss and confusion between appearance and limbs. This indicates that generating images from unaligned inputs is a highly challenging task without global features. In contrast, in B2, once the global features of the target image obtained from the prior conditional diffusion model are added, the generated person images basically match the target pose in structure. Although B2’s capability to generate images with consistent appearance is somewhat limited, it has already achieved a striking resemblance to the actual image. This shows that the refining conditional diffusion model can establish a dense correspondence between the source and target images, enhancing the contextual features. In addition, from a visual perception perspective, B3 is superior to B1 and B2 regarding detail texture, while it is slightly inferior to B2 regarding pose coordination. Finally, when we use the PCDMs of the three stages in the last column, it is visually superior to B1, B2, and B3. This indicates that when dealing with complex poses, the PCDMs of the three stages can gradually produce visually more satisfying results.
@@
To more comprehensively validate the effectiveness of the proposed refining conditional diffusion model, we apply it to other state-of-the-art (SOTA) methods. As shown in Figure 9, the first and fourth rows denote the source image and ground truth (target image). The second and third rows present the results before and after refinement via refining the conditional diffusion model. We can observe that the refining conditional diffusion model significantly improves the results of all state-of-the-art methods. For instance, when dealing with ADGAN and PIDM, our method helps to fill in minor details, such as missing belts and hats, thereby enhancing the completeness of the generated images. For methods like GFLA, DPTN, CASD, and NTED, our model can finely process textures, maintain shape and texture consistency, and generate more explicit and realistic images. These results indicate that the refining conditional diffusion model has universality across different state-of-the-art methods, offering potential improvements for various person image generation tasks.
@@ 4.3 Application
In Table 2, we further evaluate the applicability of images generated by PCDMs in a downstream task, i.e., person re-identification (Ye et al., 2021; Shen et al., 2023a). We conduct these re-identification experiments on the Market-1501 dataset, adhering to the PIDM protocol. Initially, we randomly select subsets of 20%, 40%, 60%, and 80% from the total real training set of the Market-1501 dataset, ensuring each identity is represented by at least one image. This selection process yields a new dataset. Following this, we employ BoT (Luo et al., 2019) as the base network and conduct baseline training with each subset of the data. We then generate synthetic images from the new dataset, randomly selecting those that share the same identity and pose. This synthesized data is then merged with the original dataset to train BoT. The Rank1 results are presented in Table 2. The results indicate that PCDMs significantly boost the re-identification performance compared to the baseline. Furthermore, when compared with state-of-the-art methods such as PTN, GFLA, DPTN, and PIDM, PCDMs consistently demonstrate superior performance in re-identification tasks.
@@ 5 Conclusion
This paper has demonstrated the significant potential of Progressive Conditional Diffusion Models (PCDMs) in addressing the challenges of pose-guided person image synthesis through a three-stage process. In the first stage, a simple prior conditional diffusion model is designed to predict the global features of the target image by mining the global alignment relationship between pose coordinates and image appearance. The second stage establishes a dense correspondence between the source and target images using the global features from the previous stage, and an inpainting conditional diffusion model is proposed to further align and enhance the contextual features, generating a coarse-grained person image. In the final stage, a refining conditional diffusion model is proposed to utilize the coarsely generated image from the previous stage as a condition, achieving texture restoration and enhancing fine-detail consistency. The three stages of the PCDMs work progressively to generate the final high-quality and high-fidelity synthesized image. Both qualitative and quantitative results have demonstrated the consistency and photorealism of our proposed PCDMs under challenging scenarios.
Future Work. Our method significantly improves person image synthesis quality, while the use of two additional prior and refining models leads to increased computational resource consumption and longer inference time. Future work should explore efficient methods that provide equivalent or superior quality while reducing computational overhead and inference time.